# Is ChatGPT Transforming Academics’ Writing Style?

### v1

@@ 4.3 Impact estimation and bias analysis
$\begin{split}L_{j,t}(\eta_{j})&=\frac{1}{n_{j}}\sum_{i\in I_{j}}(h_{ij}(t)-\eta_{j}(t)x_{ij}(t))^{2}\\ &=\frac{1}{n_{j}}\sum_{i\in I_{j}}(g_{ij}(t)+\xi_{ij}(t))^{2}\,.\end{split}$ (18)
If we ignored the dependency of $g_{ij}(t)$ and $\xi_{ij}(t)$ on $\eta_{j}(t)$, the estimate of ChatGPT impact would simply be given by Ordinary Least Squares (OLS) as
@@
$\begin{split}&L_{j,t,g}(\eta_{j})\\ =&\frac{1}{n_{j}}\sum_{i\in I_{j}}(h_{ij}(t)-\eta_{j}(t)x_{ij}(t)-g_{ij}(t))^{2}\\ =&\frac{1}{n_{j}}\sum_{i\in I_{j}}\xi^{2}_{ij}(t)\,.\end{split}$ (22)
@@
$\begin{split}&\frac{\partial L_{j,t,g}(\eta_{j})}{\partial\eta_{j}}\\ =&\frac{2}{n_{j}}\sum_{i\in I_{j}}\left(\eta_{j}(t)x^{2}_{ij}(t)-h_{ij}(t)x_{ij}(t)\right)\\ &-\frac{2}{n_{j}}\sum_{i\in I_{j}}\frac{\partial g_{ij}(t)}{\partial\eta_{j}(t)}\left(h_{ij}(t)-\eta_{j}(t)x_{ij}(t)-g_{ij}(t)\right)\\ &+\frac{2}{n_{j}}\sum_{i\in I_{j}}x_{ij}(t)g_{ij}(t)\,\end{split}$
@@
$\begin{split}&(\hat{\eta}^{g}_{j}(t)-\hat{\eta}_{j}(t))\sum_{i\in I_{j}}x^{2}_{ij}(t)\\ =&\sum_{i\in I_{j}}\frac{\partial g_{ij}(t)}{\partial\eta_{j}(t)}\left(h_{ij}(t)-\eta_{j}(t)x_{ij}(t)\right)\\ &-\sum_{i\in I_{j}}x_{ij}(t)g_{ij}(t)-\sum_{i\in I_{j}}g_{ij}(t)\frac{\partial g_{ij}(t)}{\partial\eta_{j}(t)}\,.\end{split}$ (23)
@@
$g_{ij}(t)=$ $\frac{\eta_{j}(t)f^{*}_{ij}(t)\epsilon_{ij}(1)}{\sqrt{\eta_{j}(t)f^{*}_{ij}(t)+\delta_{ij}(*)}}$ (25)
$-\frac{\eta_{j}(t)f^{*}_{ij}(t)\epsilon^{s}_{ij}(1)}{\sqrt{q^{d}_{ij}}}\,.$ (26)
$\begin{split}&g_{ij}(t)\frac{\partial g_{ij}(t)}{\partial\eta_{j}(t)}\\ =&g_{ij}(t)\frac{f^{*}_{ij}(t)(\eta_{j}(t)f^{*}_{ij}(t)+2\delta_{ij}(*))\epsilon_{ij}(1)}{2(\eta_{j}(t)f^{*}_{ij}(t)+\delta_{ij}(*))^{\frac{3}{2}}}\\ &-g_{ij}(t)\frac{f^{*}_{ij}\epsilon^{s}_{ij}(1)}{\sqrt{q^{d}_{ij}}}\,.\end{split}$ (27)
Removing the items with zero means, we get
$\begin{split}&\mathrm{E}\left[g_{ij}(t)\frac{\partial g_{ij}(t)}{\partial\eta_{j}(t)}\right]\\ =&\frac{\eta_{j}(t)(f^{*}_{ij}(t))^{2}(\eta_{j}(t)f^{*}_{ij}(t)+2\delta_{ij}(*))\sigma^{2}_{ij,\epsilon}}{2(\eta_{j}(t)f^{*}_{ij}(t)+\delta_{ij}(*))^{2}}\\ &+\frac{\eta_{j}(t)(f^{*}_{ij}(t))^{2}\sigma^{2}_{ij,\epsilon}}{q^{d}_{ij}}\,.\end{split}$ (28)
And the bias part is expressed as
@@
Some insights can be gained from the results above. As by definition $\eta_{j}(t)\geq 0$, the estimate $\hat{\eta}_{j}(t)$ given by Eq. (19) tends to be biased high in our model. The value of $\hat{r}_{ij}$ plays a role in the minimization of bias, as it only appears in the denominator in Eq. (29). Similarly, if the value of $\hat{r}_{ij}$ is similar for different words, then larger values of $q^{d}_{ij}$ and $f^{*}_{ij}$ will reduce the bias, as seen from Eq. (LABEL:eq:eta_bias_e_expression) – therefore, we should consider including preferentially in our analysis words with larger values of $q^{d}_{ij}$, $f^{*}_{ij}$ and $\hat{r}_{ij}$. Considering that the value of $\eta_{j}(t)$ affects the bias as well, which is not simply linear, we are led to consider adaptive or iterative criteria for word choice, which will in general depend on the true (and unknown) value of $\eta_{j}(t)$.
Case 2: Gaussian distribution for $\delta_{ij}(f_{ij})$, e.g., $\delta_{ij}(f_{ij})\sim\mathcal{N}(0,f_{ij}\sigma_{ij}^{2})$, which is justified empirically in the Appendix, Figure 14 and Figure 16. As a result,
$\begin{split}\xi_{ij}(t)=&(\hat{r}_{ij}+\epsilon^{\eta}_{ij}(q,f,t))\delta_{ij}(\eta_{j}(t)f^{*}_{ij}(t))\\ &+\delta^{\prime}_{ij}(f^{*}_{ij}(t))\\ =&\sqrt{\eta_{j}(t)f^{*}_{ij}(t)}(\hat{r}_{ij}+\epsilon^{\eta}_{ij}(q,f,t))\delta_{ij}(1)\\ &+\sqrt{f^{*}_{ij}(t)}\delta^{\prime}_{ij}(1)\end{split}$ (30)
Therefore, we can define $g^{c}_{ij}(t)$ and $\xi^{c}_{ij}(t)$:
$\begin{split}g^{c}_{ij}(t)=&\eta_{j}(t)f^{*}_{ij}(t)\epsilon^{\eta}_{ij}(q,f,t)\\ &+\sqrt{\eta_{j}(t)f^{*}_{ij}(t)}(\hat{r}_{ij}+\epsilon^{\eta}_{ij}(q,f,t))\delta_{ij}(1)\end{split}$ (31)
$\xi^{c}_{ij}(t)=$ $\sqrt{f^{*}_{ij}(t)}\delta^{\prime}_{ij}(1)$ (32)
$\begin{split}&L^{c}_{j,t,g}(\eta_{j})\\ =&\frac{1}{n_{j}}\sum_{i\in I_{j}}(h_{ij}(t)-\eta_{j}(t)x_{ij}(t)-g^{c}_{ij}(t))^{2}\\ =&\frac{1}{n_{j}}\sum_{i\in I_{j}}(\xi^{c}_{ij}(t))^{2}\,.\end{split}$ (33)
And we will get a complex expression for the bias part like Eq. (23), which gives us similar conclusions. (Some calculations are in the Appendix.)
Finding criteria for selecting the words that are included in the frequency change analysis greatly reduces the computational complexity compared to trying different word combinations. Our analysis of noise models gives some insights into these criteria, such as $q^{d}_{ij}$ and $\hat{r}_{ij}$.
4.4 Approximations in real data
Unfortunately, we cannot know the true value of $f^{*}_{ij}(t)$ in the ChatGPT era, but we can replace it with the estimation $\hat{f}^{*}_{ij}(t)$ based on the word frequency before ChatGPT was introduced. As our objective is to identify the words that ChatGPT “likes” (or “dislikes”) to use compared to academic researchers on average, we assume that the frequencies of these words should remain stable without ChatGPT, i.e., we take the average of the pre-ChatGPT periods before $T_{0}$ as following:
Considering that the noise in real data is likely highly complex, we did not estimate the variance of $\epsilon_{ij}(\cdot)$. Instead, we used ChatGPT to process additional abstracts (on top of those used to estimate $r_{ij}$), and used the resulting frequencies as calibration for the bias and noise.
@@ 4.5 Calibration and test
In order to verify the theoretical and practical validity of our approach, we used calibrations and tests, with ChatGPT-processed abstracts mixed with real abstracts.
As previous analyses have demonstrated, with the goal of reducing bias in estimation, different selected words are likely to correspond to the different (unknown) ground truth value of $\eta_{j}(t)$. Therefore, we construct $N$ different sets of abstract data for calibration, $D_{n}$, with its correspond mixed ratio of ChatGPT-processed abstracts, $\eta_{n}$, as
$(D_{n},\eta_{n}),n\in\{1,2,\dots,N\}$ (35)
and similarly for test data $T_{n^{\prime}}$ and $\eta^{\prime}_{n^{\prime}}$,
$(T_{n}^{\prime},\eta_{n}^{\prime}),n^{\prime}\in\{1,2,\dots,N^{\prime}\}.$ (36)
And for one pair of $(D_{n},\eta_{n})$ and a specific word choice requirement $q_{k}$ (for example, $r^{*}_{ij}>0.1$ and $\frac{\hat{r}_{ij}+1}{\hat{r}^{2}_{ij}}<\frac{0.1+1}{0.1^{2}}$), the efficiency can be defined as
@@
where $\hat{\eta}_{n}(D_{n},q_{k})$ is the estimate of $\eta_{n}$ using Eq. (19) and the words set $I_{j}$ can be derived from $q_{k}$, denoted $I_{j}(q_{k})$
For a given set of $q_{k}$ (examples can be found in the Appendix), we are looking for the best one minimizing $e(D_{n},\eta_{n},q_{k})$, denoted $q(D_{n},\eta_{n})$, which is the calibration part.
For the test data $T_{n^{\prime}}$, the estimate of $\eta_{n^{\prime}}$ is calculated from Eq. (19) with different $I_{j}$, based on different $q(D_{n},\eta_{n})$ obtained in the calibration procedure.
@@ 5.1 Calibration and test results
We take $\{\eta_{n}\}=\{0,0.05,1,\dots,0.45,0.5\}$ and $m=1$, which means $N=\#\{(D_{n},\eta_{n})\}=11$. Then the 11 $I_{j}$ (with possible repetitions), obtained from mixed data with 11 corresponding $\eta_{n}$ of period 12, were used for $\eta_{n}^{\prime}$ estimation in the test data (period 14). Other parameters can be found in the Appendix.
@@
The results using the same prompt for generating calibration and test data are shown in Figure 6, with injected mixed ratio (i.e., ChatGPT impact) $\eta_{n}^{\prime}$ from 0 to 0.5. It is clear that when the calibration and test sets are mixed in the same ratio, word combinations that achieve better estimates on the calibration set generally work better on the test set, as well.
Unlike in Figure 6 where we normalized the word frequency by the total number of abstracts, we normalized it by the total number of words for one period in Figure 7. The trends remain similar, albeit different in detail.
Figure 7: Similar to Figure 6, but normalized to the total number of words, rather than the total number of abstracts.
Because one may use a wide variety of prompts in practical applications, we also evaluated the robustness of our approach by adopting a different prompt for generating the test data than the one we used for calibration. The corresponding results in Figure 8 use the following prompt:
”Please rewrite the following paragraph from an academic paper:”
In this example, we add the word “please” and make it clear that this comes from an “academic paper”, replacing “revise” with “rewrite”.
Figure 8: Similar to Figure 7, but with a different prompt for test data than used in calibration data.
Although the quantitative results of our tests were not as good as before, the errors were still small at lower mixed ratios, which also illustrates the robustness of our method. This is understandable because in data generated with different prompts, not all of our previous assumptions hold, and the estimate of $\hat{r}_{ij}$ on $r_{ij}$ in our model may be biased. We can also note that most of our estimates in Figure 8 are on the high side relative to the ground truth, most likely because we use a more precise prompt for the test data here, making the frequency change rate of the relevant words higher.
@@ 5.2 Estimation from real data
The estimates of ChatGPT impact on the real data are shown in Figure 9 and Figure 10. Based on our calibration results, we chose 11 words set $I_{j}$ for different injected values of $\eta_{n}$. According to the results of the first estimation about $\eta_{j}(t)$, we found the three values of $\eta_{n}$ that were closest to the mean of the first estimation and used their optimal word set $I_{j}$ in the calibration procedure for a second estimation, leading to the triangle points shown in the figures.
@@
Figure 10: The same estimates as in Figure 9, but word frequencies were normalized by the number of words.
@@
It is important to note that our ChatGPT impact here is a relative value that corresponds to the change in word frequency from the use of simple prompts. More precise prompts, both in reality and in simulation, could potentially lead to an impact value greater than 1.
@@ 6 Conclusions
We have demonstrated here that a simple statistical analysis of word frequency is sufficient to detect and analyze the impact of ChatGPT in arXiv abstracts, which is easily extendable to other subjects and to the complete text of articles. We found convincing evidence of a change in word frequency after ChatGPT’s release, consistent with predictions obtained from simulating ChatGPT’s impact from possible users’ prompts. The most enthusiastic community (among the four we investigated) in terms of ChatGPT adoption appears to be that of computer scientists, a result that is perhaps unsurprising. Mathematicians, by contrast, are the least keen.
Our estimates are founded on a population level and based on the output of simple prompts. Using more precise prompts, it is entirely possible to achieve abstracts that are more ChatGPT-like texts than our simulations. In addition, in the real world people might use large language models other than ChatGPT to revise articles, which may have similar but not identical word preferences to ChatGPT.
Another central take-away from this article is that we can monitor the impact of ChaGPT on academic writing by using simple and transparent statistical methods (e.g., word frequencies) rather than black-box GPT detectors. At the same time, we also believe that better estimates can be made by more rigorous analysis, such as considering more complex noise terms.

### v2

@@ 4.3 Impact estimation and bias analysis
$L_{j,t}(\eta_{j})=\frac{1}{n_{j}}\sum_{i\in I_{j}}(h_{ij}(t)-\eta_{j}(t)x_{ij}(t))^{2}=\frac{1}{n_{j}}\sum_{i\in I_{j}}(g_{ij}(t)+\xi_{ij}(t))^{2}\,.$ (17)
If we ignored the dependency of $g_{ij}(t)$ and $\xi_{ij}(t)$ on $\eta_{j}(t)$, the estimate of LLM impact would simply be given by Ordinary Least Squares (OLS) as
@@
$L_{j,t,g}(\eta_{j})=\frac{1}{n_{j}}\sum_{i\in I_{j}}(h_{ij}(t)-\eta_{j}(t)x_{ij}(t)-g_{ij}(t))^{2}\\ =\frac{1}{n_{j}}\sum_{i\in I_{j}}\xi^{2}_{ij}(t)\,.$ (21)
@@
$\begin{split}\frac{\partial L_{j,t,g}(\eta_{j})}{\partial\eta_{j}}=&\frac{2}{n_{j}}\sum_{i\in I_{j}}\left(\eta_{j}(t)x^{2}_{ij}(t)-h_{ij}(t)x_{ij}(t)\right)+\frac{2}{n_{j}}\sum_{i\in I_{j}}x_{ij}(t)g_{ij}(t)\\ &-\frac{2}{n_{j}}\sum_{i\in I_{j}}\frac{\partial g_{ij}(t)}{\partial\eta_{j}(t)}\left(h_{ij}(t)-\eta_{j}(t)x_{ij}(t)-g_{ij}(t)\right)\,\end{split}$ (22)
@@
$\begin{split}(\hat{\eta}^{g}_{j}(t)-\hat{\eta}_{j}(t))\sum_{i\in I_{j}}x^{2}_{ij}(t)=&\sum_{i\in I_{j}}\frac{\partial g_{ij}(t)}{\partial\eta_{j}(t)}\left(h_{ij}(t)-\eta_{j}(t)x_{ij}(t)\right)\\ &-\sum_{i\in I_{j}}x_{ij}(t)g_{ij}(t)-\sum_{i\in I_{j}}g_{ij}(t)\frac{\partial g_{ij}(t)}{\partial\eta_{j}(t)}\,.\end{split}$ (23)
@@
$g_{ij}(t)=$ $\frac{\eta_{j}(t)f^{*}_{ij}(t)\epsilon_{ij}(1)}{\sqrt{\eta_{j}(t)f^{*}_{ij}(t)+\delta_{ij}(*)}}-\frac{\eta_{j}(t)f^{*}_{ij}(t)\epsilon^{s}_{ij}(1)}{\sqrt{q^{d}_{ij}}}\,.$ (25)
After calculations (see appendix), the bias part is expressed as
@@
Some insights can be gained from the results above. As by definition $\eta_{j}(t)\geq 0$, the estimate $\hat{\eta}_{j}(t)$ given by Eq. (18) tends to be biased high in our model. The value of $\hat{r}_{ij}$ plays a role in the minimization of bias, as it only appears in the denominator in Eq. (26).
Similarly, if the value of $\hat{r}_{ij}$ is similar for different words, then larger values of $q^{d}_{ij}$ and $f^{*}_{ij}$ will reduce the bias, as seen from Eq. (35) – therefore, we should consider including preferentially in our analysis words with larger values of $q^{d}_{ij}$, $f^{*}_{ij}$ and $|\hat{r}_{ij}|$. Considering that the value of $\eta_{j}(t)$ affects the bias as well, which is not simply linear, we are led to consider adaptive or iterative criteria for word choice, which will in general depend on the true (and unknown) value of $\eta_{j}(t)$.
Case 2: Gaussian distribution for $\delta_{ij}(f_{ij})$, e.g., $\delta_{ij}(f_{ij})\sim\mathcal{N}(0,f_{ij}\sigma_{ij}^{2})$, which is inspired by central limit theorem and justified empirically in the Appendix, Figure 8. As a result,
$\begin{split}\xi_{ij}(t)=&(\hat{r}_{ij}+\epsilon^{\eta}_{ij}(q,f,t))\delta_{ij}(\eta_{j}(t)f^{*}_{ij}(t))+\delta^{\prime}_{ij}(f^{*}_{ij}(t))\\ =&\sqrt{\eta_{j}(t)f^{*}_{ij}(t)}(\hat{r}_{ij}+\epsilon^{\eta}_{ij}(q,f,t))\delta_{ij}(1)+\sqrt{f^{*}_{ij}(t)}\delta^{\prime}_{ij}(1)\end{split}$ (27)
which gives us similar conclusions. (Some calculations are in the Appendix.)
Finding criteria for selecting the words that are included in the frequency change analysis greatly reduces the computational complexity compared to trying different word combinations. If all combinations of $n$ words are tried, that complexity rises to $O(2^{n})$. When we use word choice criteria to select several groups of words, the complexity is reduced to $O(1)$. Our analysis of noise models gives some insights into these criteria, such as $q^{d}_{ij}$ and $\hat{r}_{ij}$.
@@ 4.4 Calibration and test
In order to verify the theoretical and practical validity of our approach, we used calibrations and tests, with ChatGPT-processed abstracts mixed with real abstracts. Considering that the noise in real data is likely highly complex, we did not estimate the variance of $\epsilon_{ij}(\cdot)$. Instead, we used ChatGPT to process additional abstracts (on top of those used to estimate $r_{ij}$), and used the resulting frequencies as calibration for the bias and noise.
As previous analyses have demonstrated, with the goal of reducing bias in estimation, different selected words are likely to correspond to the different (unknown) ground truth value of $\eta_{j}(t)$. Therefore, we construct $N$ different sets of abstract data for calibration and test, $D_{n}$ and $T_{n^{\prime}}$, with its correspond mixed ratio of ChatGPT-processed abstracts, $\eta_{n}$ and $\eta^{\prime}_{n^{\prime}}$, as
$(D_{n},\eta_{n}),n\in\{1,2,\dots,N\};\quad(T_{n}^{\prime},\eta_{n}^{\prime}),n^{\prime}\in\{1,2,\dots,N^{\prime}\}.$ (28)
And for one pair of $(D_{n},\eta_{n})$ and a specific word choice requirement $q_{k}$ (for example, $q^{d}_{ij}>0.1$ and $\frac{\hat{r}_{ij}+1}{\hat{r}^{2}_{ij}}<\frac{0.1+1}{0.1^{2}}$), the efficiency can be defined as
@@
where $\hat{\eta}_{n}(D_{n},q_{k})$ is the estimate of $\eta_{n}$ using Eq. (18) and the words set $I_{j}$ can be derived from $q_{k}$, denoted $I_{j}(q_{k})$.
For a given set of $q_{k}$ (examples can be found in the Appendix), we are looking for the best one minimizing $e(D_{n},\eta_{n},q_{k})$, denoted $q(D_{n},\eta_{n})$, which is the calibration part. For the test data $T_{n^{\prime}}$, the estimate of $\eta_{n^{\prime}}$ is calculated from Eq. (18) with different $I_{j}$, based on different $q(D_{n},\eta_{n})$ obtained in the calibration procedure.
@@ 5.1 Calibration and test results
We take $\{\eta_{n}\}=\{0,0.05,0.1,\dots,0.45,0.5\}$ and $m=1$, which means $N=\#\{(D_{n},\eta_{n})\}=11$. Then the 11 $I_{j}$ (with possible repetitions), obtained from mixed data with 11 corresponding $\eta_{n}$ of period 12, were used for $\eta_{n}^{\prime}$ estimation in the test data (period 14). Other parameters can be found in the Appendix.
The results using the same prompt for generating calibration and test data are shown in Figure 2(a), with injected mixed ratio (i.e., ChatGPT impact) $\eta_{n}^{\prime}$ from 0 to 0.5. It is clear that when the calibration and test sets are mixed in the same ratio, word combinations that achieve better estimates on the calibration set generally work better on the test set, as well.
(a) Normalized to the total number of abstracts.
(b) Normalized to the total number of words.
(c) Different prompt for test data than used in calibration data.
@@
Unlike in Figure 2(a) where we normalized the word frequency by the total number of abstracts, we normalized it by the total number of words for one period in Figure 2(b). The trends remain similar, albeit different in detail.
Because one may use a wide variety of prompts in practical applications, we also evaluated the robustness of our approach by adopting a different prompt for generating the test data than the one we used for calibration. The corresponding results in Figure 2(c) use the following prompt:
"Please rewrite the following paragraph from an academic paper:"
In this example, we add the word “please” and make it clear that this comes from an “academic paper”, replacing “revise” with “rewrite”. Although the quantitative results of our tests were not as good as before, the errors were still small at lower mixed ratios, which also illustrates the robustness of our method. This is understandable because in data generated with different prompts, not all of our previous assumptions hold, and the estimate of $\hat{r}_{ij}$ on $r_{ij}$ in our model may be biased. We can also note that most of our estimates in Figure 2(c) are on the high side relative to the ground truth, most likely because we use a more precise prompt for the test data here, making the frequency change rate of the relevant words higher.
@@ 5.2 Estimation from real data
The estimates of ChatGPT impact on the real data are shown in Figure 3(a) and Figure 3(b).
(a) Normalized to the total number of abstracts.
(b) Normalized to the total number of words.
@@
Based on our calibration results, we chose 11 words set $I_{j}$ for different injected values of $\eta_{n}$. According to the results of the first estimation about $\eta_{j}(t)$, we found the three values of $\eta_{n}$ that were closest to the mean of the first estimation and used their optimal word set $I_{j}$ in the calibration procedure for a second estimation, leading to the triangle points shown in the figures.
@@
It is important to note that our ChatGPT impact or LLM impact here is a relative value that corresponds to the change in word frequency from the use of simple prompts. More precise prompts, both in reality and in simulation, could potentially lead to an impact value greater than 1.
@@ 6 Conclusions
We have demonstrated here that we can monitor the impact of LLMs in arXiv abstracts by using simple and transparent statistical methods (e.g., word frequencies), which is easily extendable to other subjects and to the complete text of articles.
Some formulas above look complicated, but with the help of the calibration, the final estimates are linear regressions, i.e., Eq. (18). And those formulas and proofs tell us which words should be theoretically selected for estimation, which to our knowledge other articles haven’t done. In addition, we also propose adaptive word selection methods that are operationally simple.
Our estimates are founded on a population level and based on the output of simple prompts. Using more precise prompts, it is entirely possible to achieve abstracts that are more ChatGPT-like (or LLM-like) texts than our simulations. In addition, in the real world people might use LLMs other than ChatGPT to revise articles, which may have similar but not identical word preferences to ChatGPT.
We found convincing evidence of a change in word frequency after ChatGPT’s release, consistent with predictions obtained from simulating LLM impact from possible users’ prompts. The most enthusiastic community (among the four we investigated) in terms of LLM adoption appears to be that of computer scientists, a result that is perhaps unsurprising. Mathematicians, by contrast, are the least keen.
Our paper illustrates the importance of words chosen. Different types of articles with different LLM impact need to be estimated using the corresponding words, which we proved theoretically under certain assumptions and verified with simulated data. Not only did we focus on words that were increasing in frequency, but we also took words that were decreasing in frequency, which are not covered in other papers.